OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 10 = x2 + 3x? A. (−2, 0) and (−5, 0) B. (2, 0) and (5, 0) C. (−2, 0) and (5, 0) D. (2, 0) and (−5, 0)

OpenStudy (hoblos):

to find x-intercept set y=0 and solve for x to find y-intercept set x=0 and solve for y

OpenStudy (experimentx):

d

OpenStudy (anonymous):

To find x-intercept put y = 0 y + 10 = x^2 + 3x 0 + 10 = x^2 + 3x x^2 + 3x - 10 = 0 x^2 + 5x - 2x - 10 =0 x(x + 5) - 2(x + 5) = 0 (x + 5)(x - 2) = 0 x + 5 = 0 , x - 2 = 0 x = -5 , x = 2 therefore, (-5, 0)(2, 0) For more help log on www.teachon.tk